Questions on Lecture and D2L Readings

For the next three questions, consider that 97 people in a community already have a job and in addition

* Joe has no job but is looking.

* Mary has a part-time job but can barely pay her bills. She wants a full-time job.

* Jim has given up looking for work and lives off his spouse’s income.

9. The labor force participation rate is

a. 1/3

b. 2/3

c. 97/99

d 99/100

10. The unemployment rate is

a. 1/3

b. 2/3

c. 1/99

d. 1/100

11. Who could be called hidden unemployed?

a. Joe, Mary, Jim

b. Mary & Jim

c. Joe only

d. Jim only
Answer #1

Q9
Answer
Option d
labor force participation rate=labor force /working age population
labor force =employed+ part time employed +looking for job
=97+1+1=99
working age population =labor force +discuraged worker(Jim)
=99+1
=100
labor force participation rate=99/100
====
Q2
Answer
option c

unemployment rate =unemployed /labor force
unemployed =lokking for job =Joe=1
Unemployement rate =1/99
=====
Q3
Hidden unemployment is the discuraged unemployment as it is not included in the unemploment but it is unemployed.
Only Jim
Option d